OpenStudy (anonymous):

how do you solve 3|2a + 7| =3a + 12

OpenStudy (harsimran_hs4):

first you need to break the modulus their will be two cases 2a + 7 >= 0 and 2a + 7 < 0 then you solve for a in both the cases and check if the solution satisfies the inequality above (case you are solving) if it satisfies then it counts as a valid solution all valid solutions is the final solution

OpenStudy (harsimran_hs4):

http://www.wolframalpha.com/input/?i=3 |2a+%2B+7|+%3D3a+%2B+12

OpenStudy (anonymous):

3|2a + 7| =3a + 12 i.e.\[ |2a + 7| =\frac{3a + 12}{3} \rightarrow |2a + 7| = a+4\] \[2a + 7 = \pm (a+4)\] \[2a + 7 = (a+4) or 2a + 7 = - (a+4)\] \[2a + 7 = a+4 or 2a + 7 = - a-4\] \[2a -a = 4-7 or 2a + a = - 7-4 \rightarrow a =-3 or 3a= -11 \] \[a =\frac{-11}{3},-3 \]

OpenStudy (anonymous):

thnk you so much :D